### v1

# The proof-theoretic strength of Constructive Second-order set theories

Abstract.
In this paper, we define constructive analogues of second-order set theories, which we will call $\mathsf{IGB}$, $\mathsf{CGB}$, $\mathsf{IKM}$, and $\mathsf{CKM}$. Each of them can be viewed as $\mathsf{IZF}$- and $\mathsf{CZF}$-analogues of Gödel-Bernays set theory $\mathsf{GB}$ and Kelley-Morse set theory $\mathsf{KM}$. We also provide their proof-theoretic strengths in terms of classical theories, and we especially prove that $\mathsf{CKM}$ and full Second-Order Arithmetic have the same proof-theoretic strength.
1. Introduction
There are two ways to formulate constructive versions of Zermelo-Frankel set theory $\mathsf{ZF}$, namely Intuitionistic Zermelo-Frankel set theory $\mathsf{IZF}$ and Constructive Zermelo-Frankel set theory $\mathsf{CZF}$. Their strengths and metamathematical properties are broadly studied by Aczel, H. Friedman, Rathjen, and others. While $\mathsf{ZF}$ is a first-order set theory–the only objects in $\mathsf{ZF}$ are sets–there is a natural demand for ‘large collections’ that are not formally sets. These objects are called classes, and second-order set theories like Gödel-Bernays set theory $\mathsf{GB}$ and Kelley-Morse set theory $\mathsf{KM}$ supports classes as objects.
It is natural to ask whether we have a constructive version of second-order set theories. The main goal of this paper is to formulate four possible second-order set theories named $\mathsf{IGB}$, $\mathsf{IKM}$, $\mathsf{CGB}$, and $\mathsf{CKM}$, and to provide their proof-theoretic strength. $\mathsf{IGB}$ and $\mathsf{CGB}$ are $\mathsf{GB}$-like second-order set theories relating to $\mathsf{IZF}$ and $\mathsf{CZF}$, respectively. (See Section 2 for a precise definition.) Since $\mathsf{GB}$ is conservative over $\mathsf{ZF}$, we may expect the same between $\mathsf{IGB}$ and $\mathsf{IZF}$, and $\mathsf{CGB}$ and $\mathsf{CZF}$ respectively. It turns out that a proof-theoretic argument for the conservativity of $\mathsf{GB}$ over $\mathsf{ZF}$ carries over (cf. [26], Problem §16.11), so that we obtain the following result:
[…]
$\mathsf{IGB}$ is conservative over $\mathsf{IZF}$ and $\mathsf{CGB}$ is conservative over $\mathsf{CZF}$. As a corollary, $\mathsf{IGB}$ and $\mathsf{IZF}$ have the same proof-theoretic strength, and the same holds for $\mathsf{CGB}$ and $\mathsf{CZF}$. ∎
Notice that $\mathsf{KM}$ is also a second-order extension of $\mathsf{ZF}$, as it is obtained from $\mathsf{GB}$ by adding class comprehension rules for arbitrary second-order formulas. $\mathsf{KM}$ is a non-conservative extension of $\mathsf{ZF}$ since $\mathsf{KM}$ proves the consistency of $\mathsf{ZF}$. We can expect the same for the constructive analogues $\mathsf{KM}$, in other words, $\mathsf{IKM}$ has a higher proof-theoretic strength than $\mathsf{IZF}$, and the same between $\mathsf{CKM}$ and $\mathsf{CZF}$. We can see that the double-negation translation between $\mathsf{IZF}$ and $\mathsf{ZF}$ can carry over a second-order setting. We may directly work with H. Friedman’s two-stage double-negation translation via non-extensional set theories, or forcing with the double-negation topology adopted in [14]. The same machinery works between $\mathsf{IKM}$ and $\mathsf{KM}$, which shows the proof-theoretic strength of $\mathsf{IKM}$:
[…]
The proof-theoretic strength of $\mathsf{IKM}$ is that of $\mathsf{KM}$. ∎
What can we say about the proof-strength of $\mathsf{CKM}$? It obviously lies between $\mathsf{CZF}$ and $\mathsf{KM}$, but even the gap between $\mathsf{CZF}$ and $\mathsf{ZF}$ is very large: It is known by [20] that the proof-theoretic strength of $\mathsf{CZF}$ is equal to that of $\mathsf{KP}$, and $\mathsf{ZF}$ proves not only the consistency of $\mathsf{KP}$ but its strengthenings like $\mathsf{KP}$ with large countable ordinals like admissible ordinals. The main result in this paper is the optimal proof-theoretic strength of $\mathsf{CKM}$ in terms of well-known classical theories:
Theorem.
The proof-theoretic strength of $\mathsf{CKM}$ is equal to that of full Second-Order Arithmetic.
Its proof is quite similar to Lubarsky’s proof [16] of the equiconsistency between full Second-Order Arithmetic and $\mathsf{CZF}$ with Full Separation.
Before starting the main content, we will briefly overview the structure of the paper. In Section 2, we review a historical account and basic notions of constructive set theory, then define the constructive versions of second-order set theories. In Section 3, we introduce preliminary notions necessary to establish the proof-theoretic strength of $\mathsf{CKM}$. In Section 4, we prove that the proof-theoretic strength of $\mathsf{CKM}$ is that of full Second-Order Arithmetic.
Acknowledgement
The initial idea of this paper stemmed from the final report for the proof theory class taught by James Walsh, and the author would like to thank him for allowing the author to write the final report about constructive set theory. The author also would like to thank Robert Lubarsky for helpful comments, and Justin Moore for improving my writing.
2. Constructive set theory
In this section, let us briefly review constructive set theories, and define constructive second-order set theories. The standard references for constructive set theory would be [4] and [5], but these might be hard to read for unmotivated readers. Thus we try to include a historical account and explanations for notions appearing in $\mathsf{CZF}$, which might help the reader to work on constructive set theory.
2.1. Historical account
Various constructive systems appeared after Bishop’s paramount ‘Constructive analysis’ [7], which showed constructive mathematics is not remote from classical mathematics. Unfortunately, Bishop formulated his concepts informally and he did not present any foundational system for his analysis, although he remained some remnants of some criteria that his system must satisfy. For example, he mentioned in his book [7] that the Axiom of Choice follows from ‘by the very meaning of the existence.’ Many consequent attempts at formulating a foundation for Bishop’s system appeared, including H. Friedman’s Intuitionistic Zermelo-Frankel set theory, Feferman’s Explicit Mathematics, Myhill’s $\mathsf{CST}$, or Martin-Löf’s type theory.
Since the set theory successfully settled the foundation crisis of mathematics, it would be natural to consider the constructive analogue of set theory as a foundation. The standard set theory for classical mathematics is Zermelo-Frankel set theory $\mathsf{ZF}$, so the constructive analogue of set theory should resemble $\mathsf{ZF}$. But classical $\mathsf{ZF}$ is deemed to be impredicative due to unrestricted quantification in the axiom of Separation and Powerset, which allows us to construct objects of a certain rank by referring to objects of higher ranks.111[17] and [8] claimed Powerset is impredicative because it does not give any rule to generate elements of a power set of a given set, and it instead assumes the universe of all sets and collects all possible subsets of the given set. See [8] for a more discussion about the impredicativity of Separation and Powerset. However, going back to Bishop’s constructive analysis that was an initial motivation for constructive foundations, Petrakis ([19], Chapter 1) stated that “… Bishop elaborated a version of dependent type theory with one universe, in order to formalise $\mathsf{BISH}$.” That is, Bishop tried to formulate his constructive analysis over a type theory, which is deemed as predicative, so a set-theoretic foundation for constructive analysis should be predicative. Then what is the constructive predicative set theory resembling $\mathsf{ZF}$? Myhill [17] and H. Friedman [10] studied a constructive version of set theory now called Intuitionistic Zermelo-Frankel set theory $\mathsf{IZF}$. It turns out by H. Friedman [11] that the proof-theoretic strength of $\mathsf{IZF}$ is equal to that of $\mathsf{ZF}$, meaning that $\mathsf{IZF}$ is highly impredicative. This rules out $\mathsf{IZF}$ as a predicative foundation for constructive analysis. Myhill [18] defined a version of ‘predicative’ constructive set theory named $\mathsf{CST}$ as an attempt to provide a foundation for Bishop’s constructive analysis [7]. However, Myhill’s formulation is not quite close to that of $\mathsf{ZF}$: for example, $\mathsf{CST}$ contains atoms for natural numbers. Years later, Aczel formulated Constructive Zermelo-Frankel set theory $\mathsf{CZF}$ and its type-theoretic interpretation in Martin-Löf type theory by modifying the type-theoretic interpretation of ordinals that appeared in the thesis of his doctoral student Leversha [15]. $\mathsf{CZF}$ is more $\mathsf{ZF}$-like than $\mathsf{CST}$, and it can be interpreted in (Martin-Löf’s) type theory (See [1], [2], [3], or [28]), so $\mathsf{CZF}$ has an reduction to a predicative theory. Therefore, $\mathsf{CZF}$ satisfies the mentioned criteria for the set-theoretic foundation for constructive analysis. Although $\mathsf{CZF}$ itself is not compatible with Choice since Choice over $\mathsf{CZF}$ implies a form of Excluded Middle, $\mathsf{CZF}$ is still compatible with weaker forms of Choice like Countable Choice or Dependent Choice that are known to be enough to develop the foundations of real analysis. (See [2] or [28] for a proof for the compatibility.) For a detailed discussion of historical aspects of constructive set theories, the reader might refer to [8].
2.2. First-order set theory
Let us start with the definition of $\mathsf{IZF}$ that is closer to that of $\mathsf{ZF}$ than $\mathsf{CZF}$:
[…]
Unlike $\mathsf{ZF}$, $\mathsf{IZF}$ uses Collection instead of Replacement as an axiom because there seems no good way to derive the full strength of $\mathsf{ZF}$ from $\mathsf{IZF}_{R}$, $\mathsf{IZF}$ with Replacement instead of Collection. In fact, H. Friedman and Ščedrov proved in [12] that $\mathsf{IZF}$ has more provably recursive functions than $\mathsf{IZF}_{R}$. Thus $\mathsf{IZF}_{R}$ is ‘weaker’ than $\mathsf{IZF}$ in some sense, although it is open whether $\mathsf{IZF}$ proves the consistency of $\mathsf{IZF}_{R}$.
$\mathsf{CZF}$ appeared first in Aczel’s paper [1] about the type-theoretic interpretation of set theory.
[…]
  • $\Delta_{0}$-Separation: $\{x\in a\mid\phi(x,p)\}$ exists if $\phi$ is a bounded formula.
  • Strong Collection: If $\forall x\in a\exists y\phi(x,y)$, then there is $b$ such that
    $\forall x\in a\exists y\in b\phi(x,y)\text{ and }\forall y\in b\exists x\in a\phi(x,y).$
  • Subset Collection: For any $a$ and $b$, we can find $c$ such that for any $u$, if $\forall x\in a\exists y\in b\phi(x,y,u)$, then we can find $d\in c$ such that
    $\forall x\in a\exists y\in d\phi(x,y,u)\text{ and }\forall y\in d\exists x\in a\phi(x,y,u).$
The formulation of $\mathsf{CZF}$ may appear to be a bit strange since the choice of axioms is less motivated (especially, Subset Collection). The following notions could be helpful in approaching these two axioms:
[…]
Let $A$ and $B$ be classes.
  • A relation $R\subseteq A\times B$ is a multi-valued function if $\forall x\in A\exists y\langle x,y\rangle\in R$. In this case, we use the notation $R\colon A\rightrightarrows B$ and we call $A$ a domain of $R$.
  • If $R\colon A\rightrightarrows B$ further satisfies $\forall y\in B\exists x\in A\langle x,y\rangle\in R$, then we call $B$ a subimage of $R$ and use the notation $R\colon A\mathrel{\mathrlap{\leftleftarrows}{\rightrightarrows}}B$.
In the $\mathsf{CZF}$-world, multi-valued functions are more fundamental than functions and subsets. The notion of subsets is associated with power sets, which are viewed as impredicative. But we may ask why functions are not ‘fundamental’ like multi-valued functions. The reason comes from the type-theoretic interpretation of $\mathsf{CZF}$. When we interpret Strong Collection over the type interpretation, for example, we extract a choice function from the statement $\forall x\in a\exists y\phi(x,y)$. However, the choice function $f$ we get is never extensional: The choice function $f$ depends on the ‘type-theoretic expression’ of sets $x\in a$. We can still form a set-sized ‘mapping’ $G=\{\langle x,f(x)\rangle\mid x\in a\}\subseteq\{\langle x,y\rangle\mid x\in a\land\phi(x,y)\}$, but $G$ as a graph is no longer functional since a given set $x\in a$ may have different type-theoretic expressions, and $f(x)$ depends on that expression. Thus $G$ ends up being a multi-valued function. A similar situation also appears around interpreting Subset Collection.
We can see that if $R$ is functional, that is, for each $x$ there is at most one $y$ such that $\langle x,y\rangle\in R$, then a subimage must be an image of $R$. However, unlike images, subimages of a given multi-valued function need not be unique. Under this terminology, we can formulate Strong Collection and Subset Collection as variations of Replacement and Powerset in $\mathsf{ZF}$. For Strong Collection, let us recall that Replacement states
For every set $a$ and a class function $F\colon a\to V$, $F$ has an image $b$ of $F$.
By replacing ‘function’ and ‘image’ with ‘multi-valued function’ and ‘subimage,’ we get the statement of Strong Collection:
For every set $a$ and a class multi-valued function $F\colon a\rightrightarrows V$, $F$ has a subimage $b$ of $F$.
The case for Subset Collection requires more work because it requires functional formulation for Powerset. The reason we take a functional formulation instead of a usual formulation of Powerset is the axiom of Powerset itself is impredicative as Myhill [18] pointed out. On the other hand, the notion of functions is not impredicative since “a mapping or function is a rule, a finite object which can actually be given” according to Myhill ([18], p351).
[…]
Powerset is equivalent to the following claim: for given sets $a$ and $b$ we can find a set $c$ such that if $f\colon a\to b$ is any (set or class) function, then $c$ contains the image of $f$.
Proof.
The forward direction is clear. For the other direction, let $c$ be the set containing all images of functions from $a$ to $a\times 2$. For each $u\subseteq a$, consider the function whose existence is guaranteed by Replacement:
$f_{u}(x)=\begin{cases}\langle x,1\rangle&\text{if }x\in u,\\ \langle x,0\rangle&\text{if }x\notin u.\end{cases}$
Then $\operatorname{ran}f_{u}=(a\setminus u)\times\{0\}\cup u\times\{1\}$. Hence the set
$\{\{x\in a:\langle x,1\rangle\in u\}\mid u\in c\},$
which exists by Replacement, contains all subsets of $a$. ∎
Since the set of all images of functions $f\colon a\to b$ is an example of a collection of definable class functions, we can further strengthen the previous lemma as follows:
[…]
Powerset is equivalent to the following statement: for given sets $a,b$ and a class family of classes $\{F_{u}\mid u\in V\}$ defined by $F_{u}=\{\langle x,y\rangle\mid\phi(x,y,u)\}$, we can find a set $c$ such that if $F_{u}\colon a\to b$ is a function, then $\operatorname{ran}F_{u}\in c$.
Proof.
The forward direction is clear. The converse direction follows from subsection 2.2 since we can take $\{F_{u}\mid u\in V\}=\{f\mid f\colon a\to b\}$. ∎
By modifying the above lemma, we can notably get the statement of Subset Collection:
For given sets $a,b$ and a class family of classes $\{F_{u}\mid u\in V\}$ defined by $F_{u}=\{\langle x,y\rangle\mid\phi(x,y,u)\}$, we can find a set $c$ such that if $F_{u}\colon a\rightrightarrows b$ is a multi-valued function, then $c$ contains a subimage of $F_{u}$.
One expected consequence of Subset Collection is that it implies the function set exists:
[…]
$\mathsf{CZF}$ proves for each sets $a$ and $b$, ${}^{a}b=\{f\mid f\colon a\to b\}$ exists.
Proof.
Consider the following class families $\{F_{u}\mid u\in V\}$ defined as follows:
$F_{u}=\{\langle x,\langle x,y\rangle\rangle\mid\langle x,y\rangle\in u\cap(a\times b)\}.$
We can see that if $u$ is a function from $a$ to $b$, then $F_{u}$ is a function from $a$ to $a\times b$ and the image of $F_{u}$ is $u$. Now let $c$ be a set of subimages of $\{F_{u}\mid u\in V,\,F_{u}\colon a\rightrightarrows b\}$. If $u\colon a\to b$ is a function, then it is the image of $F_{u}$, which is the unique subimage of $F_{u}$ since $F_{u}$ is a function. Hence $u\in c$. Therefore, by Bounded Separation, we have
${}^{a}b=\{u\in c\mid u\colon a\to b\}.\qed$
However, it is known that $\mathsf{CZF}$ does not prove the existence of a powerset: Working over $\mathsf{CZF}$ with Powerset, we can construct $V_{\omega+\omega}$, which is a model of Intuitionistic Zermelo set theory $\mathsf{IZ}$. Furthermore, [13] showed that $\mathsf{IZ}$ interprets classical Zermelo set theory by Friedman’s double-negation interpretation. On the other hand, [20] proved that the proof-theoretic strength of $\mathsf{CZF}$ is equal to that of $\mathsf{KP}$, which is strictly weaker than that of Zermelo set theory. Also, the proof-theoretic strength of $\mathsf{CZF}$ with Powerset in terms of classical theories is provided by Rathjen [22]: He showed that the proof-theoretic strength of $\mathsf{CZF}$ with Powerset is equal to a Power-Kripke set theory $\mathsf{KP}(\mathcal{P})$, $\mathsf{KP}$ with Separation and Collection for $\Delta_{0}^{\mathcal{P}}$-formulas, which is strictly stronger than Zermelo set theory. For the precise definition of terminologies, see [22].
2.3. Second-order set theory
We will define two types of second-order set theories as we classically did. There are two second-order extensions of $\mathsf{ZF}$, namely Gödel-Bernays set theory $\mathsf{GB}$ and Kelly-Morse set theory $\mathsf{KM}$. Some literatures may use different names (like $\mathsf{NBG}$ for $\mathsf{GB}$, $\mathsf{MK}$ for $\mathsf{KM}$) but we follow terminologies in [29]. The main difference between $\mathsf{GB}$ and $\mathsf{KM}$ is while $\mathsf{GB}$ cannot use second-order formulas to formulate new classes, $\mathsf{KM}$ can.
The following definition defines the constructive analogue of $\mathsf{GB}$:
[…]
Second-order set theories are defined over a two-sorted language, each will mean sets and classes. We use $\forall^{0}$ and $\exists^{0}$ to quantify sets, and $\forall^{1}$ and $\exists^{1}$ to quantify classes.
Axioms of $\mathsf{CGB}$ are first-order axioms of $\mathsf{CZF}$ with the following second-order axioms:
  • Class Extensionality: two classes are equal if they have the same set members. Formally,
    $\forall^{1}X,Y[X=Y\leftrightarrow\forall^{0}x(x\in X\leftrightarrow x\in Y)].$
  • Elementary Comprehension: if $\phi(x,p,C)$ is a first-order formula with a class parameter $C$, then there is a class $A$ such that $A=\{x\mid\phi(x,p,C)\}$. Formally,
    $\forall^{0}p\forall^{1}C\exists^{1}A\forall^{0}x[x\in A\leftrightarrow\phi(x,p,C)].$
  • Class Set Induction: if $A$ is a class, and if we know a set $x$ is a member of $A$ if every element of $x$ belongs to $A$, then $A$ is the class of all sets. Formally,
    $\forall^{1}A\big{[}[\forall^{0}x(\forall^{0}y\in x(y\in A)\to x\in A)]\to\forall^{0}x(x\in A)\big{]}.$
  • Class Strong Collection: if $R$ is a class multi-valued function from a set $a$ to the class of all sets, then there is a set $b$ which is an ‘image’ of $a$ under $R$. Formally,
    $\forall^{1}R\forall^{0}a[R\colon a\rightrightarrows V\to\exists^{0}b(R\colon a\mathrel{\mathrlap{\leftleftarrows}{\rightrightarrows}}b)].$
Axioms of $\mathsf{IGB}$ are obtained from $\mathsf{CGB}$ by adding axioms of $\mathsf{IZF}$ with the following axiom:
  • Class Separation: for any class $A$ and a set $a$, $A\cap a$ is a set. Formally,
    $\forall^{1}A\forall^{0}a\exists^{0}b\forall^{0}x[x\in b\leftrightarrow(x\in A\land x\in a)]$
and putting Class Collection instead of Class Strong Collection:
  • Class Collection: if $R$ is a class multi-valued function from a set $a$ to the class of all sets, then there is a set $b$ such that $a$ ‘projects’ to $b$ by $R$. Formally,
    $\forall^{1}R\forall^{0}a[R\colon a\rightrightarrows V\to\exists^{0}b(R\colon a\rightrightarrows b)].$
The reader should be cautioned that Elementary Comprehension does not tell anything about forming sets. That is, $\{x\in a\mid\phi(x)\}$ is a class in general unless $\phi$ is $\Delta_{0}$. Also, as $\mathsf{IZF}$ implies Strong Collection, $\mathsf{IGB}$ also proves Class Strong Collection from Class Collection:
[…]
$\mathsf{IGB}$ proves Class Strong Collection.
Proof.
It follows the usual proof for Strong Collection over $\mathsf{IZF}$: Suppose that $R\colon a\rightrightarrows V$ is a multi-valued function and let $b$ be a set such that $R\colon a\rightrightarrows b$. Then take $c=\{y\in b\mid\exists x\in a(\langle x,y\rangle\in R)\}$, which exists by Class Separation. Then $R\colon a\mathrel{\mathrlap{\leftleftarrows}{\rightrightarrows}}c$. ∎
The constructive analogues $\mathsf{CKM}$ and $\mathsf{IKM}$ of $\mathsf{KM}$ are defined from $\mathsf{CGB}$ and $\mathsf{IGB}$ by adding a stronger class formation rule:
[…]
$\mathsf{CKM}$ and $\mathsf{IKM}$ are obtained from $\mathsf{CGB}$ and $\mathsf{IGB}$ by adding the following axiom:
  • Full Comprehension: if $\phi(x,p,C)$ is a second-order formula with a class parameter $C$, then there is a class $A$ such that $A=\{x\mid\phi(x,p,C)\}$.
Also, the following alternative form of Class Collection eliminates the need to define ordered pairs for its formulation. It is useful when we work with formulas directly, so we want to avoid additional complexity introduced by the definition of ordered pairs.
[…]
Class Collection is equivalent to the following claim: for any formula $\phi(x,y,p,A)$ with no class quantifiers,
$\forall^{1}A\forall^{0}p\forall^{0}a[(\forall^{0}x\in a\exists^{0}y\phi(x,y,p,A))\to\exists^{0}b(\forall^{0}x\in a\exists^{0}y\in b\phi(x,y,p,A))].$
Proof.
For the left-to-right, consider $R=\{\langle x,y\rangle\mid x\in a\land\phi(x,y,p,A)\}$. Then $R\colon a\rightrightarrows V$, then apply the Class Collection. For the right-to-left, observe that the formula $\langle x,y\rangle\in R$ is a formula without class quantifiers. ∎
Similarly, we can formulate the following form of Strong Class Collection over $\mathsf{CGB}$:
[…]
Strong Class Collection is equivalent to the following claim: for any formula $\phi(x,y,p,A)$ with no class quantifiers,
$\forall^{1}A\forall^{0}p\forall^{0}a[(\forall^{0}x\in a\exists^{0}y\phi(x,y,p,A))\to\exists^{0}b(\forall^{0}x\in a\exists^{0}y\in b\phi(x,y,p,A)\land\forall^{0}y\in b\exists^{0}x\in a\phi(x,y,p,A))].\qed$
3. Other preliminary notions
In the remaining part of the paper, we will provide a way to reduce $\mathsf{CKM}$ into Second-Order Arithmetic. The main idea is tweaking Lubarsky’s interpretation defined in [16]. The main idea of Lubarsky’s proof is coding sets as well-founded trees combined with Kleene’s first realizability algebra. We will not review Lubarsky’s proof in full detail, but we still review the necessary details in his paper.
3.1. Second-Order Arithmetic
In this paper, we work over Second-Order Arithmetic unless specified. But what is Second-Order Arithmetic? Usually, we characterize Second-Order Arithmetic with first-order Peano arithmetic with Induction and Comprehension for any second-order formulas. In this paper, we additionally assume choice axioms over Second-Order Arithmetic. Before defining the choice schemes, let us mention that we denote $(Y)_{i}=\{m\mid(i,m)\in Y\}$ to mention a coded family of sets.
[…]
$\mathsf{\Sigma^{1}_{\infty}-AC}$ is the following assertion for any second-order formula $\phi(n,X)$:
$\forall n\exists X\ \phi(n,X)\to\exists Z\forall n\phi(n,(Z)_{n}).$
$\mathsf{\Sigma^{1}_{\infty}-DC}$ is the following assertion for each second-order formula $\phi(n,X,Y)$:
$\forall n\forall X\exists Y\ \phi(n,X,Y)\to\exists Z\forall n\ \phi(n,(Z)^{n},(Z)_{n}),$
where $(Z)^{n}=\{(i,m)\in Z\mid i<n\}$
It is known that adding $\mathsf{\Sigma^{1}_{\infty}-AC}$ and $\mathsf{\Sigma^{1}_{\infty}-DC}$ to full Second-Order Arithmetic does not change its proof-theoretic strength. For its proof, see Theorem VII.6.16 of [23]. On the other hand, [9] showed that $\mathsf{\Sigma^{1}_{\infty}-AC}$ does not suffice to prove a fragment of $\mathsf{\Sigma^{1}_{\infty}-DC}$.
3.2. Finite sequences and trees
It is known that we can code a finite sequence of natural numbers into a single natural number, and we denote a code of the sequence $a_{0},a_{1},\cdots,a_{n}$ as $\vec{a}=\langle a_{0},a_{1},\cdots,a_{n}\rangle$. For a finite sequence $\vec{a}=\langle a_{0},a_{1},\cdots,a_{n}\rangle$, we denote its length as $\operatorname{len}(\vec{a})$, and its restriction $\langle a_{0},\cdots,a_{k-1}\rangle$ as $\vec{a}\mkern-5.0mu\upharpoonright k$. For two finite sequences $\vec{a}$ and $\vec{b}$, we denote their concatenation as $\vec{a}^{\frown}\vec{b}$, whose length is $\operatorname{len}\vec{a}+\operatorname{len}\vec{b}$.
Now we can code trees in terms of sets of finite sequences with certain conditions:
[…]
A set $S$ is a tree if it satisfies the following conditions:
  1. (1)
    $\langle\rangle\in S$,
  2. (2)
    If $\vec{a}\in S$ and $k\leq\operatorname{len}(\vec{a})$, then $\vec{a}\mkern-5.0mu\upharpoonright k\in S$.
[…]
$\operatorname{immd}(S)=\{a\mid\langle a\rangle\in S\}.$
A tree $S$ is well-founded if it has no infinite path over there, that is,
$\lnot\exists f\forall n[\langle f(0),\cdots,f(n)\rangle\in S].$
We denote that tree $S$ is well-founded as $\operatorname{WF}(S)$. It is known that full Second-Order Arithmetic with the help of $\mathsf{\Sigma^{1}_{\infty}-DC}$ satisfies $\mathsf{\Pi^{1}_{\infty}-TI}$, which is equivalent over full Second-Order Arithmetic to the assertion that all well-founded relations satisfy transfinite induction for arbitrary second-order formulas.
[…]
Second-Order Arithmetic proves the following: If $\prec$ is a well-founded relation over the domain $D$, and if $\phi(x)$ is a second-order formula, then the following holds:
$\forall x\in D[(\forall y\prec x\phi(y))\to\phi(x)]\to\forall x\in D\ \phi(x)\qed$
Now let $S$ be a well-founded tree and for $\vec{a},\vec{b}\in S$, define $\vec{a}\prec_{S}\vec{b}$ if $\vec{a}$ is an immediate successor of $\vec{b}$, that is, $\exists n(\vec{a}=\vec{b}^{\frown}\langle n\rangle)$. Then $\prec_{S}$ forms a well-founded relation over $S$, so we have the following:
[…]
Suppose that $S$ is a well-founded tree and $\phi(x)$ is a second-order statement. If we have
$\forall\vec{a}\in S[\forall\vec{a}^{\frown}\langle n\rangle\in S\ \phi(\vec{a}^{\frown}\langle n\rangle)]\to\phi(\vec{a}),$
then we have $\forall\vec{a}\in S\phi(\vec{a})$. ∎
Also, we can prove the following simultaneous induction over two well-founded trees:
[…]
Suppose that $S$ and $T$ are well-founded trees and $\phi(x,y)$ is a second-order statement. If we have
$\forall\vec{a}\in S\forall\vec{b}\in T[\forall\vec{a}^{\frown}\langle n\rangle\in S\forall\vec{b}^{\frown}\langle m\rangle\in T\ \phi(\vec{a}^{\frown}\langle n\rangle,\vec{b}^{\frown}\langle m\rangle)]\to\phi(\vec{a},\vec{b}),$
then we have $\forall\vec{a}\in S\forall\vec{b}\in T\phi(\vec{a},\vec{b})$.
Proof.
Consider the relation pairwise comparison relation $\prec_{S}\times\prec_{T}$ given by
$(p,q)\mathrel{(\prec_{S}\times\prec_{T})}(r,s)\iff(p\prec_{S}r)\land(q\prec_{T}s)$
over $S\times T$. Then we can see that $\prec_{S}\times\prec_{T}$ is a well-founded relation over $S\times T$. Thus the desired conclusion follows from subsection 3.2. ∎
3.3. Kleene’s first combinatory algebra
Structures known as partial combinatory algebras or applicative structures provide a way to handle generalized computable functions in an algebraic manner. We define partial combinatory algebras in an axiomatic manner:
[…]
$\mathsf{PCA}$ is the intuitionistic first-order theory defined over the language with two constant symbols $\mathbf{k}$ and $\mathbf{s}$, with a ternary relational symbol $\mathsf{App}$. We denote $\mathsf{App}(t_{0},t_{1},t_{2})$ as $t_{0}t_{1}\simeq t_{2}$.
In practice, we represent statements over $\mathsf{PCA}$ in terms of partial terms with the symbol $\simeq$. Partial terms are (non-associative) juxtapositions of terms over $\mathsf{PCA}$, and here $t_{0}t_{1}\cdots t_{n}$ means $(\cdots(t_{0}t_{1})t_{2}\cdots)$.
We also use the following notations: If $s$ and $t$ are partial terms and $t$ is not a variable, then define
$s\simeq t\iff\forall y(s\simeq y\leftrightarrow t\simeq y).$
If $a$ is a free variable, define
$s\simeq a\iff\begin{cases}s=a&\text{If $s$ is a term of $\mathsf{PCA}$,}\\ \exists x\exists y[s_{0}\simeq x\land s_{1}\simeq y\land xy\simeq a]&\text{if $s$ is of the form $(s_{0}s_{1})$.}\end{cases}$
Finally, $t\downarrow$ means $\exists y(t\simeq y)$. Now define the axioms of $\mathsf{PCA}$ as follows:
  1. (1)
    $ab\simeq c_{0}$ and $ab\simeq c_{1}$ implies $c_{0}=c_{1}$,
  2. (2)
    $(\mathbf{k}ab)\downarrow$ and $\mathbf{k}ab\simeq a$,
  3. (3)
    $(\mathbf{s}ab)\downarrow$ and $\mathbf{s}abc\simeq ac(bc)$.
The structure of $\mathsf{PCA}$ looks simple, but it is known that $\mathsf{PCA}$ is strong enough to interpret lambda calculus, and enough to prove the recursion theorem:
[…]
Let $t$ be a partial term and $x$ be a variable. Then we can construct a term $\lambda x.t$ whose free variables are that of $t$ except for $x$, such that $\mathsf{PCA}$ proves $\lambda x.t\downarrow$ and $(\lambda x.t)u\simeq t[x/u]$ for any partial terms $u$. ∎
[…]
We can find a partial term $\mathbf{r}$ such that $\mathsf{PCA}$ proves $\mathbf{r}x\downarrow$ and $\mathbf{r}xy\simeq x(\mathbf{r}x)y$. ∎
It is known that Turing machines form a partial combinatory algebra, also known as Kleene’s first algebra. In general, for a given set $X$, the collection of all $X$-computable functions $\{e\}^{X}$ form a partial combinatory algebra.
[…]
For a set $X\subseteq\mathbb{N}$ and $a,b,c\in\mathbb{N}$, let us define $ab\simeq c$ if $\{a\}^{X}(b)=c$. Then we get a partial combinatory algebra if we appropriately interpret $\mathbf{s}$ and $\mathbf{k}$. Furthermore, we can interpret $\mathbf{s}$ and $\mathbf{k}$ uniform to $X$. ∎
It is known that we can conservatively extend $\mathsf{PCA}$ so that it includes terms for pairing functions, projections, and natural numbers. However, we only work with Kleene’s first algebra, so such an extension in full generality is unnecessary for our purpose. Despite that, we should introduce some notation under the virtue of this extension for later use:
[…]
We write $e^{X}a$ instead of $\{e\}^{X}(a)$ if the notation does not cause any confusion. Especially, if $X=\emptyset$, we write $ea$ instead of $e^{\emptyset}a$. Also, we use $\mathbf{p}xy$, $(x)_{0}$, $(x)_{1}$ to denote a pairing function and the corresponding projection functions.
For more details on $\mathsf{PCA}$, see [27] or [6].
3.4. Hyperarithmetical sets
A set $X\subseteq\mathbb{N}$ is arithmetical if it is defined by a first-order formula, or equivalently, its Turing degree is below $\emptyset^{(n)}$ for some $n$. We define hyperarithmetical sets as a generalization of an arithmetical set as follows:
[…]
A set $X$ is hyperarithmetical if the Turing degree of $X$ is below $\emptyset^{(\alpha)}$ for some recursive ordinal $\alpha$.
However, the above definition is not directly formalizable within Second-Order Arithmetic. Thus let us take a bypass to formulate within Second-Order Arithmetic as [23] did.
[…]
Fix a set $X\subseteq\mathbb{N}$. Define $\mathcal{O}_{+}^{X}$ as the set of all $(e,i)$, where $e$ is an $X$-recursive index for a $X$-recursive linear ordering and $i\in^{X}_{e})$, where $\leq^{X}_{e}$ is the decoded linear order from $e$ and $X$. For $a,b\in\mathcal{O}_{+}^{X}$, define $b<^{X}_{\mathcal{O}}a$ if $a=(e,i)$, $b=(e,j)$ for some $e$, $i$, $j$, and $j<^{X}_{e}i$ holds.
For $a\in\mathcal{O}_{+}^{X}$, we say $a\in\mathcal{O}^{X}$ if there is no $f\colon\mathbb{N}\to\mathcal{O}_{+}^{X}$ such that $f(0)=a$ and $f(n+1)<^{X}_{\mathcal{O}}f(n)$ for all $n\in\mathbb{N}$.
[…]
For a set $X$, define $X^{\prime}=\{e\mid\{e\}^{X}(e)\downarrow\}$. Also, $C(a,X,Y)$ is the following statement:
$C(a,X,Y)=\{(0,m)\mid m\in X\}\cup\{(b+1,m)\mid b<^{X}_{\mathcal{O}}a\land m\in(Y)_{b+1}\},$
take $\mathrm{H}(a,X,Y)$ by $\mathcal{O}^{X}_{+}(a)\land Y=C(a,X,Y)\land[\forall b<^{X}_{\mathcal{O}}a\ (Y)_{b+1}=C(b,X,Y)^{\prime}]$.
Intuitively, $\mathcal{O}^{X}$ is the set of $X$-recursive ordinal representations, and $C(a,X,Y)$ is a code of an $a$-indexed family starting from $X$ and the next stages are given by $Y$. $\mathrm{H}(a,X,Y)$ means $Y$ codes an $a$-indexed family of iterated Turing jumps, starting from $X$ and for each $b<^{X}_{\mathcal{O}}a$, the $b$-th set is the Turing jump of the code of the family of previous Turing jumps of $X$. It is known that a weaker subsystem of full Second-Order Arithmetic named $\mathsf{ATR}_{0}$ proves an iterated Turing jump always exists:
[…]
$\mathsf{ATR}_{0}$ proves if $a\in\mathcal{O}^{X}$, then $\forall X\exists Y\ \mathrm{H}(a,X,Y)$. Furthermore, such $Y$ is always uniquely determined by $a$ and $X$, which will be called $X^{(a)}$.∎
Now we can define hyperarithmetical sets as follows:
[…]
A set $Y$ is hyperarithmetical in $X$ if there is $a\in\mathcal{O}^{X}$ such that $Y\leq_{T}X^{(a)}$. $Y$ is hyperarithmetical if it is hyperarithmetical in the empty set $\emptyset$. A statement $\phi(n)$ is hyperarithmetical if there is a hyperarithmetical set $Y$ such that $\phi(n)$ is equivalent to $n\in Y$.
Note that being hyperarithmetical is equivalent to being $\Delta^{1}_{1}$. For more details on hyperarithmetical sets, see [23].
4. The Proof-theoretic strength of $\mathsf{CKM}$
Before defining the realizability model, let us define which subsets of $\mathbb{N}$ will be interpreted as sets and classes of $\mathsf{CKM}$. The basic idea is viewing sets as well-founded trees as [16] did, but we further restrict that the trees are hyperarithmetical. This restriction was motivated by the fact that $L_{\omega_{1}^{CK}}$ is a model of $\mathsf{KP}$ with the proof-theoretic association between $\mathsf{KP}$ and $\mathsf{CZF}$, and that a subset of $\mathbb{N}$ is hyperarithmetical exactly when it is a member of $L_{\omega_{1}^{CK}}$.
Throughout this section, we work over full Second-Order Arithmetic with $\mathsf{\Sigma^{1}_{\infty}-AC}_{0}$.
[…]
Let $S$ be a tree. Then we say $\operatorname{Set}(S)$ holds if $S$ is well-founded and for each $\vec{a}\in S$, $\operatorname{immd}(\vec{a}^{S})$ is hyperarithmetical. $\operatorname{Class}(S)$ holds if for all $n\in\operatorname{immd}(S)$, $\operatorname{Set}(n^{S})$ holds. We say $S$ is a set name or a class name if it satisfies $\operatorname{Set}(S)$ or $\operatorname{Class}(S)$ respectively.
Note that $S$ itself need not be hyperarithmetical even when $\operatorname{Set}(S)$ holds. Also, the following is immediate:
[…]
If $\operatorname{Set}(S)$ and $\vec{a}\in S$, then $\operatorname{Set}(\vec{a}^{S})$. In particular if $\operatorname{Class}(S)$, $\vec{a}\neq\langle\rangle$, and $\vec{a}\in S$, then $\operatorname{Set}(\vec{a}^{S})$. ∎
Now we define the realizability relation recursively as follows:
[…]
We distinguish bounded quantifiers and unbounded quantifiers as syntactically distinct objects, and bounded quantifiers must be bounded by a set name and not by a class name. We always assume that $S$ and $T$ in this definition always set names or class names. Also, we distinguish set quantifiers and class quantifiers by using the superscript 0 or 1. Let us define $e\Vdash\phi$ inductively as follows:
  • $e\Vdash S\in T$ if and only if $(e)_{0}\in\operatorname{immd}(T)$ and $(e)_{1}\Vdash S=(e)_{0}^{T}$.
  • $e\Vdash S=T$ if and only if for $a\in\operatorname{immd}(S)$, $(e)_{0}a\Vdash a^{S}\in T$ and for $b\in\operatorname{immd}(T)$, $(e)_{1}b\Vdash b^{T}\in S$.
  • $e\Vdash\phi\land\psi$ if and only if $(e)_{0}\Vdash\phi$ and $(e)_{1}\Vdash\psi$.
  • $e\Vdash\phi\lor\psi$ if and only if either $(e)_{0}=0$ and $(e)_{1}\Vdash\phi$, or $(e)_{0}=1$ and $(e)_{1}\Vdash\psi$.
  • $e\Vdash\phi\to\psi$ if and only if for all $f$, $f\Vdash\phi$ implies $ef\Vdash\psi$.
  • $e\Vdash\forall X\in S\phi(S)$ if and only if for all $a\in\operatorname{immd}(S)$, $ea\Vdash\phi(a^{S})$.
  • $e\Vdash\exists X\in S\phi(S)$ if and only if $(e)_{0}\in\operatorname{immd}(S)$ and $(e)_{1}\Vdash\phi((e)_{0}^{S})$.
  • $e\Vdash\forall^{0}X\phi(X)$ if and only $\forall X[\operatorname{Set}(X)\to e\Vdash\phi(X)]$.
  • $e\Vdash\exists^{0}X\phi(X)$ if and only $\exists X[\operatorname{Set}(X)\land e\Vdash\phi(X)]$.
  • $e\Vdash\forall^{1}X\phi(X)$ if and only $\forall X[\operatorname{Class}(X)\to e\Vdash\phi(X)]$.
  • $e\Vdash\exists^{1}X\phi(X)$ if and only $\exists X[\operatorname{Class}(X)\land e\Vdash\phi(X)]$.
We claim that we can realize all axioms of $\mathsf{CKM}$. Before going to the main proof, let us state some lemmas that have a pivotal role in our main argument.
[…]
Let $S$ and $T$ be set names. Then the statement $e\Vdash S=T$ is a hyperarithmetical statement for $e$.
Proof.
By unpacking the definition, we can see that $e\Vdash S=T$ is the conjunction of the following two statements:
  • For all $a\in\operatorname{immd}(S)$, $(e)_{0}a\downarrow$, $((e)_{0}a)_{0}\in\operatorname{immd}(T)$, and $((e)_{0}a)_{1}\Vdash a^{S}=((e)_{0}a)_{0}^{T}$, and
  • For all $b\in\operatorname{immd}(T)$, $(e)_{1}b\downarrow$, $((e)_{1}b)_{0}\in\operatorname{immd}(S)$, and $((e)_{1}b)_{1}\Vdash b^{T}=((e)_{1}b)_{0}^{S}$.
We appeal to subsection 3.2 to show $e\Vdash S=T$ is hyperarithmetical. That is, we inductively prove that for each $\vec{a}\in S$ and $\vec{b}\in T$, the statement $e\Vdash\vec{a}^{S}=\vec{b}^{T}$ is hyperarithmetical to $e$. By replacing $\vec{a}^{S}$ and $\vec{b}^{T}$ to $S$ and $T$ respectively, it suffices to see that $e\Vdash S=T$ is hyperarithmetical if for each $a\in\operatorname{immd}(S)$ and $b\in\operatorname{immd}(T)$, $f\Vdash a^{S}=b^{T}$ is hyperarithmetical statement for $f$.
Under the given hypothesis, we can see that each of $((e)_{0}a)_{1}\Vdash a^{S}=((e)_{0}a)_{0}^{T}$ and $((e)_{1}b)_{1}\Vdash b^{T}=((e)_{1}b)_{0}^{S}$ are hyperarithmetical statements. Furthermore, the statements $(e)_{0}a\downarrow$ and $(e)_{1}b$ are arithmetical, and we assumed that $\operatorname{immd}(S)$ and $\operatorname{immd}(T)$ are hyperarithmetical. Hence we can see that the unpacked statement of $e\Vdash S=T$ is hyperarithmetical. ∎
[…]
Let $\phi$ be a bounded formula of the language of set theory. Then $e\Vdash\phi$ is hyperarithmetical.
Proof.
We use induction on $\phi$. Suppose $S$ and $T$ are set names. The case $e\Vdash S=T$ follows from Section 4, and the case $e\Vdash S\in T$ is an immediate corollary of it. The remaining cases are also direct since adding propositional connectives or bounded quantifiers only increases the complexity of $e\Vdash\phi$ at most the Turing jump of the sum of realizing sentences of previous ones. ∎
[…]
All valid statements of intuitionistic first-order logic are realized. Furthermore, bounded quantifiers are correct in the sense that we can realize $\forall X\in S\phi(X)\leftrightarrow\forall X(X\in S\to\phi(X))$ and $\exists X\in S\phi(X)\leftrightarrow\exists X(X\in S\land\phi(X))$.
Proof.
Its proof is easy but repetitive and tedious. Its proof is quite similar to proofs in earlier literature like [21], [24], or [25], so we omit it. ∎
Theorem 4.1.
All axioms of $\mathsf{CKM}$ are realized.
Proof.
Most parts of the proof in [16] carry over. Especially, the proof for Extensionality, Set Induction, Pairing, Union, and Infinity in [16] works directly to our setting. Hence we only state proofs that cannot be directly applied to our setting.
  1. (1)
    Bounded Separation: Suppose that $S$ is a set name. For a bounded formula $\phi(X,Y)$ and a set name $T$, consider
    $S_{T,\phi}=\{\langle\rangle\}\cup\{(\mathbf{p}fn)^{\frown}\vec{a}\mid n^{\frown}\vec{a}\in S\land f\Vdash\phi(n^{S},T)\}.$
    By Section 4 with that $S$ is a set name, $f\Vdash\phi(n^{S},T)$ is hyperarithmetical to $f$ and $n$. It is clear that $S_{T,\phi}$ is well-founded, so $S_{T,\phi}$ is a set name.
  2. (2)
    Class Comprehension: The same proof for Bounded Separation works. The difference is that $T$ can be a class name and $\phi$ need not be bounded.
  3. (3)
    Second-order Strong Collection: The same proof in [16] works, but let us state it again for clarity. Suppose that $e\Vdash\forall X\in S\exists Y\phi(X,Y)$. Then we have
    $\forall a\in\operatorname{immd}(S)\exists Y\ \operatorname{Set}(Y)\land[ea\Vdash\phi(a^{S},Y)].$
    Now apply $\mathsf{\Sigma^{1}_{\infty}-AC}$ to get a set $Z\subseteq\mathbb{N}$ such that
    $\forall a\in\operatorname{immd}(S)\ \operatorname{Set}((Z)_{a})\land[ea\Vdash\phi(a^{S},(Z)_{a})].$
    Now set $T=\{\langle\rangle\}\cup\{\langle a\rangle^{\frown}\vec{b}\mid a\in\operatorname{immd}(S)\land\vec{b}\in(Z)_{a}\}$. Here $T$ itself may not be hyperarithmetical, but $\operatorname{immd}(T)$ is hyperarithmetical. Thus we have $\operatorname{Set}(T)$ and
    $\forall a\in\operatorname{immd}(S)[ea\Vdash\phi(a^{S},a^{T})].$
    This means we have
    $\lambda a.(\mathbf{p}a(ea))\Vdash\forall X\in S\exists Y\in T\ \phi(X,Y),$
    so we can see that $\lambda e\lambda a.(\mathbf{p}a(ea))$ realizes Strong Collection for classes.
  4. (4)
    Class Subset Collection: [16] gave a proof for realizing Subset Collection, but let us rework with the proof presented in [24]. Suppose that $S$ and $T$ are set names. For each $e\in\mathbb{N}$, define
    $\overline{e}=\{\langle f\rangle^{\frown}\vec{b}\mid f\in\operatorname{immd}(S)\ \land\ (ef)\downarrow\ \land\ \langle ef\rangle^{\frown}\vec{b}\in T\}.$
    We can see that $\operatorname{immd}(\overline{e})=\operatorname{immd}(S)$ and for every $\vec{b}$ such that $\langle ef\rangle^{\frown}\vec{b}\in B$, $\operatorname{immd}(\langle f\rangle^{\frown}\vec{b}^{\overline{e}})=\operatorname{immd}(\langle ef\rangle^{\frown}\vec{b}^{T})$. Therefore $\operatorname{Set}(\overline{e})$. Note that $\overline{e}$ is defined if and only if
    $\forall f\in\operatorname{immd}(S)(ef)\downarrow\land\ (ef)\in\operatorname{immd}(T),$
    which is a hyperarithmetical formula with parameter $e$. Now let
    $D=\{\langle\rangle\}\cup\{\langle e\rangle^{\frown}\vec{a}\mid e\in\operatorname{immd}(S),\ \overline{e}\text{ is defined and }\vec{a}\in\overline{e}\}.$
    Then $D$ is also a set name and $e^{D}=\overline{e}$. Now we claim that $D$ witnesses Second-order Subset Collection.
    Suppose that $e\Vdash\forall X\in S\exists Y\in T\phi(X,Y)$, where $\phi$ is any second-order statement with parameters allowed. Let us set $e^{\prime}=\lambda x.(ex)_{0}$, then we can see that $\overline{e}^{\prime}$ is defined and so $e^{\prime}\in\operatorname{immd}(D)$.
    Now for each $f\in\operatorname{immd}(S)$, we can see that $(ef)_{0}\in\operatorname{immd}(T)$ and $(ef)_{1}\Vdash\phi(f^{S},(ef)_{0}^{T})$, and $(ef)_{0}^{T}=f^{\overline{e}^{\prime}}$. Hence we have $\lambda f.\mathbf{p}f((ef)_{1})\Vdash\forall X\in S\exists Y\in\overline{e}^{\prime}\phi(X,Y)$ and also $\lambda f.\mathbf{p}f((ef)_{1})\Vdash\forall Y\in\overline{e}^{\prime}\exists X\in S\phi(X,Y)$. In sum, we can see that the realizer
    $\lambda e.\mathbf{p}e^{\prime}[\mathbf{p}(\lambda f.\mathbf{p}f((ef)_{1}))(\lambda f.\mathbf{p}f((ef)_{1}))]$
    realizes Second-order Subset Collection. ∎
It is easy to prove that $\mathsf{CKM}$ interprets intuitionistic full Second-Order Arithmetic, by interpreting second-order predicates over $\mathbb{N}$ to subclasses of $\omega$ over $\mathsf{CKM}$. Since intuitionistic full Second-Order Arithmetic interprets classical full Second-Order Arithmetic via a double-negation translation, we have
Theorem 4.2.
The proof-theoretic strength of $\mathsf{CKM}$ is that of Second-Order Arithmetic. ∎